Three times the sum of three consecutive integers is 72. What are the integers?
Answer by rfadrogane(214)

Three times the sum of three consecutive integers is 72. What are the integers?
Sol'n:
3(x+x+1+x+2)=72
3x+3=24
3x=24-3
3x=21
x=7
x+1=8
x+2=9
thus the integers are: 7,8 & 9 ---answer
check: 3(7+8+9)=72
72=72

^_^